OpenStudy (anonymous):

Find the equation of the line having the given slope and containing the given point. m=7/8, (7,-2) The equation of the line is y=

OpenStudy (anonymous):

IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

OpenStudy (anonymous):

Which formula would it be?

OpenStudy (anonymous):

Point-Slope one!

OpenStudy (anonymous):

How would I write this in slope-intercept form?

OpenStudy (anonymous):

Do you have slope m yet?

OpenStudy (anonymous):

BTW, are you sure you read my instruction correctly?

OpenStudy (anonymous):

It told e to simplify answer and type it in slope-intercept form

OpenStudy (anonymous):

INPUT YOUR VALUES INTO THE POINTSLOPE FORM THEN SIMPLIFY AND CONVERT TO SLOPE INTERCEPT FORM

OpenStudy (anonymous):

@jmallis Can you start with plugging the values into point-slope form?

OpenStudy (anonymous):

y=7/8x+(7,2) ?

OpenStudy (anonymous):

Where do you get 7.2 ???

OpenStudy (anonymous):

(7,-2)

OpenStudy (anonymous):

Oh, (7, -2) means x = 7, y -2

OpenStudy (anonymous):

Now can you replug the values in?

OpenStudy (anonymous):

so y=7/8(7)+(-2)

OpenStudy (anonymous):

@jmallis you have huge problem with following the instruction, do you (Sighs) You can't go directly to point intercept without knowing y intercept, so you need 2 steps: first write in Point-Slope form, then simplify it into Slope-Intercept!

OpenStudy (anonymous):

Point-Slope: y +2 = (7/8) ( x - 7) Now distribute 7/8 in, simplify it into Slope-intercept form!